A satellite has a mass of 100 kg and is located at 2 X10⁶ m above the surface of the Earth.


a) What is the Potential energy associated with the satellite at this location?


b) What is the magnitude of the gravitational force on the satellite?



1
Expert's answer
2022-03-08T12:32:33-0500

a) the potential energy of the satellite-Earth system

"U=-\frac{GMM'}{r}"

Where M is mass of earth and othercof soatellite.

"U=-\frac{ 6.67×\times10^{−11} \times5.97 × 10^{24}}{2\times10^6 }"

"=" "-20\times10^{7} J"

b) magnitude of the gravitational force exerted by the earth on the satellite

"F=\frac{GMM'}{r^2}"

"=\frac{ 6.67×\times10^{−11} \times5.97 × 10^{24}}{(2\times10^6 )^2}"

"=100N"